I tried a substitution problem on my own.I am tryng my best .I do not think I did it the proper way but I think I have the answer and I would like your opinion and help if I got it wrong.Please.The problem is:
x-4y=5
x-2y=1
The answer I got is -3,-2. Is this correct?
Thankyou

Answer by checkley71(8403)

X-4Y=5
-X+2Y=-1
-----------------------------
-2Y=4
Y=4/-2
Y=-2 ANSWER.
X-4*-2=5
X+8=5
X=5-8
X=-3 (-3,-2) IS THE CORRECT ANSWER.
-(-3)+2(-2)=-1
3-4=-1
-1=-1
-----------------------------------------------------------
As I stated earlier, I thank you very much. Do you feel like helping me do another for the
road. The problem is: 5x+2y =7 x-y = 0
---------------------------------
x-y or x=y
5*y+2y=7
5y+2y=7
7y=7
y=7/7
y=1 answer.
x=1 answer.